OpenStudy (anonymous):

simplify the expression(cos^2 x + sin^2 x)/( cot^2 x- csc^2 x)

OpenStudy (anonymous):

answer is -1. shall i simplify??

OpenStudy (anonymous):

yes please and thankyou

OpenStudy (anonymous):

cos^2 x + sin^2 x=1 cot^2 x= cos^2 x/sin^2 x csc^2 x=1/sin^2 x cos^2 x-1/sin^2 x= cos^2 x - sin^2 x - cos^2 x/sin^2 x= -1 1/-1=-1 do you understand??

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

can you also help me with this problem? Use basic identities to simplify the expression. (cos^2(theta)/sin^2(theta)+ csc θ sin θ

OpenStudy (anonymous):

cos^2 θ/sin^2 θ + 1/sinθ * sinθ=cos^2 θ/sin^2 θ+1 cos^2 θ+ sin^2 θ/ sin^2 θ=1/sin^2 θ=csc^2 θ did you get it??

OpenStudy (anonymous):

yeah thanks alot

OpenStudy (anonymous):

you're welcome again :)